The net present value for Proposals X and Y has been calculated on the basis of the data given below.

Proposal X

Proposal Y

Amount to be invested

$145,000

$280,000

Total present value of net cash flow

172,000

320,000

Net present value

27,000

40,000

Determine the present value index for each proposal.
Answer #1

This is a refinement of the Net Present Value method. Present value index is yet another time-adjusted method which measures the present value of returns per rupee invested. This index is also called as Benefit-cost ratio or profitability index. It may be defined as the ratio which is obtained by dividing the PV of the future cash inflows by the present value of cash outflows.

It can be put up in the form of the following formula as follows :

Present Value index (PI) = Present value of future cash inflows
Present value of future cash outflows

Accept/Reject Criterion:

Under this method, a project would be qualified to be accepted if its PI exceeds one. If PI equals one, the firm is indifferent to the project. If PI is lesser than one, the project would be rejected. When two or more projects are compared, the project which has the highest PI will be considered. This is because, higher the PI, greater is the profitability of the project.

Profitability Index method has got similar benefits like the Net Present value method which are as follows :

  1. It considers time value of money.
  2. It takes into account the cash inflows and outflows throughout the economic life of the project.
  3. The benefits of considering cash flows rather than accounting profit also applies to this method.
  4. It is capable of handling any discount rate to determine the present value of cash flows to suit the prevailing market condition.
  5. Though PI method is almost similar to NPV method and has got the same advantages, the former is still a better measure because PI measures the relative profitability and NPV, being an absolute measure.

Calculation of Present Value Index :

As NPV = Present Value of Cash Inflow – Present Value of Cash Outflow

Present Value of future cash inflows = NPV + Present Value of Cash Outflow

For Proposal X ,

Present Value of future cash inflows = 172000   (or 27000+145000)

Present Value of future cash outflows = 145000

Present Value Index = 172000/145000 = 1.18

For Proposal Y ,

Present Value of future cash inflows = 320000 (or 40000 + 280000)

Present Value of future cash outflows = 280000

Present Value Index = 172000/145000 = 1.14

As PVI of Proposal X exceed , proposal X is more profitable.